# Charge functions for high dimensional partitions

### v1

@@ Abstract
To construct a BPS algebra with representations furnished by n-dimensional partitions, the first step is to construct the eigenvalue of the Cartan operators acting on them. The generating function of the eigenvalues are called the charge function. It has an important property that for each partition, the poles of the function correspond to the projection of the boxes which can be added to or removed from the partition legally. The charge functions of lower dimensional partitions, i.e., Young diagrams for 2D, plane partitions for 3D and solid partitions for 4D, are already given in the literature. In this paper, we propose an expression of the charge function for arbitrary odd dimensional partitions and have it proved for 5D case. Some explicit numerical tests for 7D and 9D case are also conducted to confirm our formula.
@@ 1 Introduction
This paper is organized as follows. In section 2, we review some necessary concepts about partitions in arbitrary dimension. In section 3, we conjecture an expression of the charge function for any odd dimension. Section 4 serves as a preparation for the following proof. In section 5, we demonstrate that the conjectured formula indeed satisfies the property of the charge function, as long as Lemma 5 holds. Then in section 6, we give a complete proof for 5D case. For higher odd dimensions, we offer a partial proof, and perform Monte Carlo sampling tests for 7D and 9D. We conclude in section 7. The details of the proof for Lemma 1-4 are collected in appendix A.
@@ 4 Definitions
$D_{\Delta^{(n)}}\coloneqq\left\{c=\sum_{i}l_{i}h_{i}\in\mathcal{P}\ \bigm|\ \sum_{i}l_{i}\vec{e}_{i}\in C(\Delta^{(n)})\right\}.$ (4.3)
Note that $\left|C(\Delta^{(n)})\right|=\left|D_{\Delta^{(n)}}\right|$. And we introduce another set to collect the simple pole of the partition
@@
$G(\vec{\square})=\{\Delta^{n}\in P_{n}\ |\ \exists\tilde{\vec{\square}}\in C(\Delta^{n}),\tilde{c}(\tilde{\vec{\square}})=c(\vec{\square})\}.$ (4.5)
@@
$L{(\Delta^{(n)},\vec{\square})}\coloneqq\Delta^{(n)}\cap\{\tilde{\vec{\square}}=\sum_{i=1}^{n}\tilde{l_{i}}\vec{{e_{i}}}\,|\,\tilde{l_{i}}-l_{i}\geq 0,\,\forall\tilde{l_{i}}\}$. (4.23)
@@
A $d$-dimensional hypercube $HC$ with its origin $\vec{\square}$ is defined as a subspace with $2^{d}$ elements in $\mathbb{Z}^{n}$:
@@ 6.1 Some analytically check of Lemma 5
First we give a proof of Lemma 5a:
@@
Noting that $\forall\,k\in\{3,5\dots 2k-1\}$, there’s one cluster centered at each nkn-k neighbor of $\sum_{i=1}^{d}e_{n_{i}}$ that contributes a pole to $\sum_{i=1}^{d}e_{n_{i}}$. In a word, each $n-m,(1\leq m\leq n-1)$ neighbor contributes $(-1)^{n+1}$ poles. This is the same predicted by (7.1).
@@ 6.2 Numerical proof of Lemma 5
Figure 5: Visualization of a special $n=5$ partition (200 boxes). Each subplot shows its projection onto $h_{1},h_{2},h_{3}$, with horizontal rightward as increasing $h_{4}$ and vertical downward as increasing $h_{5}$. Red/green squares denote the positions which can add new boxes or remove existing boxes (melting rule), and black dots represent simple poles of the charge function. Their perfect coincidence confirms consistent judgments between the two methods, verifying the universality of our method for general cases.
@@ 2. Pole contribution breaks sign symmetry for even $n$
Appendix A Proof of Lemma 1-4
@@ A.1 Lemma 1
which is the melting rule of $\Delta^{(n)}-L(\Delta^{(n)},\vec{\square})$ that proof the Lemma by definition. ∎

### v2

# Charge functions for odd dimensional partitions

@@ Abstract
To construct a BPS algebra with representations furnished by n-dimensional partitions, the first step is to construct the eigenvalue of the Cartan operators acting on them. The generating function of the eigenvalues is called the charge function. It has an important property that for each partition, the poles of the function correspond to the projection of the boxes which can be added to or removed from the partition legally. The charge functions of lower dimensional partitions, i.e., Young diagrams for 2D, plane partitions for 3D and solid partitions for 4D, are already given in the literature. In this paper, we propose an expression of the charge function for arbitrary odd dimensional partitions and have it proved for 5D case. Some explicit numerical tests for 7D and 9D case are also conducted to confirm our formula.
@@ 1 Introduction
This paper is organized as follows. In section 2, we review some necessary concepts about partitions in arbitrary dimension. In section 3, we conjecture an expression of the charge function for any odd dimension. Section 4 serves as a preparation for the following proof. In section 5, we demonstrate that the conjectured formula indeed satisfies the property of the charge function, as long as Lemma 5 holds. Then in section 6, we give a complete proof for 5D case. For higher odd dimensions, we offer a partial proof, and perform Monte Carlo sampling tests for 7D and 9D. We conclude in section 7. The details of the proofs for Lemma 1-4 are collected in appendix A.
@@ 4 Definitions
$D_{\Delta^{(n)}}\coloneqq\left\{c=\sum_{i}l_{i}h_{i}\in\mathcal{P}\ \bigm|\ \sum_{i}l_{i}\vec{e}_{i}\in C_{\Delta^{(n)}}\right\}.$ (4.3)
Note that $\left|C_{\Delta^{(n)}}\right|=\left|D_{\Delta^{(n)}}\right|$. And we introduce another set to collect the simple pole of the partition
@@
$G(\vec{\square})=\{\Delta^{(n)}\in P_{n}\ |\ \exists\,\tilde{\vec{\square}}\in C_{\Delta^{(n)}},\tilde{c}(\tilde{\vec{\square}})=c(\vec{\square})\}.$ (4.5)
@@
$L{(\Delta^{(n)},\vec{\square})}\coloneqq\Delta^{(n)}\cap\{\tilde{\vec{\square}}=\sum_{i=1}^{n}\tilde{l_{i}}\vec{{e_{i}}}\,|\,\tilde{l_{i}}-l_{i}\geq 0,\,\forall\,i\}.$ (4.23)
@@
A $d$-dimensional hypercube $HC$ with its origin $\vec{\square}$ is defined as a subspace with $2^{d}$ elements in $\mathbb{Z}_{\geq 0}^{n}$:
@@ 6.1 Some analytically check of Lemma 5
We give a proof of Lemma 5a:
@@
Noting that $\forall\,k\in\{3,5\dots 2K-1\}$, there’s one cluster centered at each nkn-k neighbor of $\sum_{i=1}^{d}\vec{e}_{n_{i}}$ that contributes a pole to $\sum_{i=1}^{d}\vec{e}_{n_{i}}$. In a word, each $n-m,(1\leq m\leq n-1)$ neighbor contributes $(-1)^{n+1}$ poles.
@@ 6.2 Numerical proof of Lemma 5
Figure 5: Visualization of a special $n=5$ partition (200 boxes). Each subplot shows its projection onto $\vec{e}_{1},\vec{e}_{2},\vec{e}_{3}$, with horizontal rightward as increasing $\vec{e}_{4}$ and vertical downward as increasing $\vec{e}_{5}$. Red/green squares denote the positions which can add new boxes or remove existing boxes (melting rule), and black dots represent simple poles of the charge function. Their perfect coincidence confirms consistent judgments between the two methods, verifying the universality of our method for general cases.
@@ 2. Pole contribution breaks sign symmetry for even $n$
Appendix A Proofs of Lemma 1-4
@@ A.1 Lemma 1
which is equivalent to the melting rule of $\Delta^{(n)}-L(\Delta^{(n)},\vec{\square})$, thus proving the Lemma by definition. ∎